A sinusoidal electromagnetic wave from a radio station passes perpendicularly through an open window that has area of 0.550m^2. At the window, the electric field of the wave has an rms value 1.60E-2 V/m .


How much energy does this wave carry through the window in a 30.0 s commercial?
Answer #1

Intensity I = power/area = Erms^2/uoC
where uo = constnat = 4pi*10^-7 H/m
C = velocity of light = 3e8 m/s
Emax = Erms* 1.414
Emax and Bmax are related by Emax = Bmax C
and Power = eenrgy/time

so power = 1.6*1.6 *10^-4 * 0.55/(4pi*10^-7 *3e8)

P = 3.736*10-7

eenrgy E = pt

E = 3.736*30

E = 1.1201*10^-5 J